OpenStudy (anonymous):

The perimeter of a rectangle is 80x. The base is 7 times the height. In terms of x, what are the dimensions of the rectangle?

OpenStudy (anonymous):

okay one sec i got u

OpenStudy (anonymous):

Perimeter = 2B+2H ^ B = 7H 80X = 2(7H) + 2H 80X = 16H 10X = 2H So, H = 5X and B = 35X Area of the rectangle = BH = (5X) (35X) A = 175X^2

OpenStudy (anonymous):

good looks bro

OpenStudy (anonymous):

Find the area of a trapezoid in which base1 is 9 units, base2 is 7 units, and the height is 3 units.

OpenStudy (anonymous):

dont call me bro ima chick

OpenStudy (stamp):

lol

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

area of trapezoid = [(b1+b2)h]/2 sub it in you get... area = [(12+14)13]/2 =[26x13]/2 =338/2 =169 units square

OpenStudy (anonymous):

its okay:)

OpenStudy (anonymous):

What are the dimensions of a rectangle that has a perimeter of 44 units and an area of 105 square units?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

i cant get it

OpenStudy (anonymous):

Find the value of x if the area of the trapezoid is 54 square units.

OpenStudy (anonymous):

its very easy try it. use a formula nd u will get it